OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of -7 and y-intercept 6.

jimthompson5910 (jim_thompson5910):

Remember, in general, the equation of a line is y = mx+b where m is the slope and b is the y-intercept.

OpenStudy (anonymous):

Ok so are you going to help me or not?

jimthompson5910 (jim_thompson5910):

so what is the value of m?

OpenStudy (anonymous):

I dont know what?

jimthompson5910 (jim_thompson5910):

the slope is given as -7, so m = -7 since m is the slope

OpenStudy (anonymous):

Ok

jimthompson5910 (jim_thompson5910):

what about b?

OpenStudy (anonymous):

6?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so y = mx+b turns into what???

OpenStudy (anonymous):

6= idont know the rest

jimthompson5910 (jim_thompson5910):

you agree that m = -7 and b = 6 correct?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

so we replace m with -7 and we replace b with 6 to go from y = mx+b to y = -7x+6

OpenStudy (anonymous):

Ok and? then what

jimthompson5910 (jim_thompson5910):

and that's it, that's the answer since it's in slope intercept form

OpenStudy (anonymous):

so i write online that it's y= -7x+6

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

Can you help me with another one real fast @jim_thompson5910

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

k thanks. here it is. write the equation in slope-intercept form of the line that has a slope of 5 and y-intercept of -3.

jimthompson5910 (jim_thompson5910):

Same idea, but this time the slope is 5, so m = 5, The value of b is then what?

OpenStudy (anonymous):

-3

jimthompson5910 (jim_thompson5910):

good, so b = -3

jimthompson5910 (jim_thompson5910):

So y = mx+b turns into what?

OpenStudy (anonymous):

y=5x+-3

jimthompson5910 (jim_thompson5910):

very good, we can write +-3 as just -3 So y = 5x+-3 can be written as y = 5x-3

jimthompson5910 (jim_thompson5910):

so the final answer is y = 5x-3

OpenStudy (anonymous):

So thats the final answer??!

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

Thanks

jimthompson5910 (jim_thompson5910):

you're welcome